OpenStudy (unklerhaukus):

\[y'' + x(y')^2 = 0\]

ganeshie8 (ganeshie8):

Split it as : y' = u u' + xu^2 = 0 second equation is separable

OpenStudy (unklerhaukus):

\[u'+xu^2 = 0\\ u'/u^2 = -x\\ -1/u = -(x^2+c)/2\\ 1/u = (x^2+c)/2\\ y' = 2/(x^2+c)\\ y = \int\frac2{x^2+c}\mathrm dx\]

OpenStudy (unklerhaukus):

but (2/a) arctan (x/a) + b doesn't solve the original eqation

OpenStudy (baru):

guess :p y=c

OpenStudy (unklerhaukus):

y = c is only one solution

OpenStudy (baru):

if its any help :P its definitely not anything in the form x^n

OpenStudy (unklerhaukus):

i think it's inverse hyperbolic

OpenStudy (faiqraees):

How about using a substitution like z =1/y or something?

OpenStudy (agent0smith):

Hmm why doesn't ganeshie's method work http://www.wolframalpha.com/input/?i=y%27%27+%2B+x(y%27)%5E2+%3D+0

OpenStudy (unklerhaukus):

the sign of c

OpenStudy (alekos):

You're right uncle rhaukus it's to do with the sign of c. Inverse hyper tan answer

OpenStudy (yagosik):

It does solve the original equation. \[y=\frac{ 2 }{ c _{1} }\arctan (\frac{ x }{ c _{1} })+ c _{2}\] \[y \prime=\frac{ 2 }{ 1+x^2 }\] \[y \prime \prime=\frac{ -4x }{ (1+x^2)^2 }\] \[y \prime \prime+x(y \prime)^2=-\frac{ 4x }{ (1+x^2)^2 }+x(\frac{ 2 }{ (1+x^2) })^2=0\]

OpenStudy (yagosik):

One more thing. If c in \[y=\int\limits_{}^{}\frac{ 2 }{ x^2+c }dx \] is a negative than we'll have another answer.

OpenStudy (unklerhaukus):

aha yes, it does solve the equation (when i don't forget the -ve sign on the second derivative) ,thank you @yagosik. The answer provided in the back of the book was y = c or y = - ln x + c, but y = - ln x + c y' = -1/x y'' = 1/x^2 y''+x(y')2 = 1/x^2 + x(-1/x)^2 = 1/x^2 + 1/x ≠ 0

OpenStudy (alekos):

Well done @yagosik

OpenStudy (alekos):

y=(2/c1)arctanh (x/c1) + c2 Is also an answer

OpenStudy (unklerhaukus):

How can I get the hyperbolic solution from the equation?

ganeshie8 (ganeshie8):

arctah(x) = iarctan(-ix)

ganeshie8 (ganeshie8):

\[\begin{align} y &= c_2 - \dfrac{\sqrt{2}\color{red}{arctanh (x/\sqrt{2c_1})}}{\sqrt{c_1}}\\~\\ &= c_2 - \dfrac{\sqrt{2}\color{red}{iarctan (-ix/\sqrt{2c_1})}}{\sqrt{c_1}}\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

Let \(-i/\sqrt{2c_1} = 1/c_1'\) and we should be good..

OpenStudy (alekos):

Brilliant